I need help with this word problem:
A very skinny man is standing by a tall tree.The man's shadow is 14 feet and the tree has a shadow that is 40 feet long. If he is 5.5 feet tall,find how tall the tree is.
Thank you.

Answer by checkley77(12844)

5.5/14=X/40 CROSS MULTIPLY.
14X=5.5*40
14X=220
X=220/14
X=15.71 FT. TALL FOR THE TREE.